OpenStudy (anonymous):

prove : √(2√(3√(4√(...))) < 3

OpenStudy (turingtest):

all square roots?\[\sqrt{2\sqrt{3\sqrt{4\sqrt{...}}}}<3\]

OpenStudy (anonymous):

yeah

OpenStudy (turingtest):

I'm helpless on these things, but enjoy the show :)

OpenStudy (anonymous):

:)

OpenStudy (anonymous):

we must think of something like this \[\sqrt{2\sqrt{3\sqrt{4\sqrt{...}}}}<\sqrt{2\sqrt{4\sqrt{8\sqrt{...}}}}=2^{\frac{1}{2}+\frac{1}{2}+\frac{3}{8}+... }=4\sqrt{2}\]

OpenStudy (turingtest):

can you explain how\[2^{\frac{1}{2}+\frac{1}{2}+\frac{3}{8}+... }=4\sqrt{2}\]?

OpenStudy (anonymous):

\[\sum_{n=1}^{\infty} \frac{n}{2^n}=2\]

OpenStudy (anonymous):

@KingGeorge what do u think?

OpenStudy (kinggeorge):

The only other possible method I know of would be to square both sides, solve for the square root, and find a pattern in the right side.

OpenStudy (kinggeorge):

\[\sqrt{2\sqrt{3\sqrt{4\sqrt{...}}}}\overset{?}<3\]\[\sqrt{3\sqrt{4\sqrt{...}}}\overset{?}<\frac{9}{2}<3\]\[\vdots\]

OpenStudy (kinggeorge):

I have my doubts about this way working however. @mukushla's method will probably be your best bet.

OpenStudy (anonymous):

@TuringTest ************we must think of something like this \[\sqrt{2\sqrt{3\sqrt{4\sqrt{...}}}}<\sqrt{2\sqrt{4\sqrt{8\sqrt{...}}}}=2^{\frac{1}{2}+\frac{1}{2}+\frac{3}{8}+... }=4\]

OpenStudy (anonymous):

reverse induction will work here

OpenStudy (anonymous):

yeah its working....... let \[f(n)=\sqrt{m\sqrt{(m+1)\sqrt{(m+2)\sqrt{...\sqrt{n}}}}}\] for \(m=n\) we have \(\sqrt{n}<n+1\) true... let suppose this is true for \(m+1\) it means \[\sqrt{(m+1)\sqrt{(m+2)\sqrt{(m+3)\sqrt{...\sqrt{n}}}}}<m+2\] we show that its true for \(m\) \[\sqrt{m\sqrt{(m+1)\sqrt{(m+2)\sqrt{...\sqrt{n}}}}}<\sqrt{m(m+2)}<m+1\] so the statement is true for all \(n \in N\) let \(m=2\) we are done...

OpenStudy (anonymous):

tnx @Neemo the definition of \(f\) we assume that \(m\le n\) and we want to show that for all \(m\le n\) this statement is true : \(f(n)<m+1\)

OpenStudy (anonymous):

Oke, thanks mukushla and any one help me solved this problem...

OpenStudy (anonymous):

yw :)